OpenStudy (anonymous):

Find a cartesian equation for the curve and identify its graph. (r^2)(cos(2theta))=1

OpenStudy (anonymous):

use this \[\cos 2 \theta=\cos^2 \theta-\sin^2 \theta\\\] and for Converting between polar and Cartesian coordinates \[x=r \cos \theta \ \ \ and \ \ \ y=r \sin \theta\]